OpenStudy (anonymous):

If you are finding the line perpendicular to x + 3y = 5, what is the slope of the perpendicular line?

OpenStudy (bahrom7893):

find the slope of this line first, get it into y=mx+b form

OpenStudy (bahrom7893):

x+3y=5 3y=-x+5 y=(-1/3)x+(5/3)

OpenStudy (anonymous):

y= -1/3x +5/3 so slope of perpendicular line is 3

OpenStudy (anonymous):

A. -3/5 B. 1/3 C. 3 D. -3 these are the answers given

OpenStudy (bahrom7893):

C

OpenStudy (anonymous):

Thank you guys! :D